Find the real root of the equation 𝑥𝑒𝑥=2 correct upto two decimal places using the method of false position.


1
Expert's answer
2022-03-08T14:04:02-0500

"f(x)=xe^x-2"


For "x-1" "f(1)=1x^1-2=0=0.718"


For "x=0.5" "f(0.5)\u00d7e^{0.5}-2=-1.17"


As the value of "f(x)" at "x=0.5" is -ve and "x=1.0" it is positive. The root lies between "x=0.5" and "x=1"


1st Step

"x_0=0.5" and "x_1=1"


"x_2=x_0-\frac{(x_1-x_0)}{f(x_1)-(x_0)}" "f(x_0)"


"=0.5-\frac{1-0.5}{0.718-(-1.17)}" "(-1.17)"


"=0.8098"

"f(x_2)=f(0.8098)=0.8098e^{0.8098}-2"

"=0.8098"


2nd Step

Root lies between "x_1=1" and "x_2=0.8098"


"x_3=x_2-\frac{x_1-x_2}{f(x_1)-f(x_2)}" "f(x_2)"


"=0
.8098-\frac{1-0.8098}{0.718+0.18}" "(-0.18)"

"=0.8479"

"f(x_2)=f(0.8479)=0.8479-e^{0\n8479}-2"


"=-0.02"


3rd Step

Next root lies between "x_3=0\n.8479" and "x_1=1"


"x_4=x_3-\frac{x_1-x_3}{f(x_1)-f(x_3)}" "f(x_3)"


"=0
.8479-\frac{1-0.8479}{0.718-(-0.02)}" "(-0.02)"

"=0.8519"


"f(x_4)=0.8519e^{0.8519}-2=-0.0031"


4th Step

Next value of "x" lies in between "x_1=1" and "x_4=0.8519"


"x_5=x_4-\frac{x_1-x_4}{f(x_1)-f(x_4)}" "f(x_4)"


"=0.8519-\frac{1-0.8519}{0.718+0.0031}" "(0.0031)=0.8525"


Approximate root to 4d.p "=0.8525"